# Error threshold in active steering protocols for few-qubit systems

### v1

@@ I Introduction
Below we explore a route to state preparation and stabilization, which builds upon the recently developed “active steering” protocols [35, 36, 37, 38, 39, 40, 41]. (We here follow Ref. [42] and use the notion “steering” as proxy for “guiding” the quantum system; this notion differs from “quantum steering” in quantum information theory [43].) In active steering protocols, one employs feedback operations based on weak measurements of the system, instead of performing projective syndrome measurements as in standard QEC. Typically, in weak measurements, the system qubits of interest are weakly coupled to detector qubits which are then projectively measured [19]. The key idea is then to perform periodically repeated, non-invasive measurements of the system state, where the detector output determines the application of subsequent operations in each time step, e.g., in order to prepare, stabilize, or manipulate a desired quantum state. This scheme should be contrasted with “passive steering” protocols [42, 44], where the weak measurement outcomes are simply discarded. In effect, passive steering protocols are equivalent to the driven-dissipative passive protection schemes mentioned above.
While active steering protocols require one to perform many measurements, which in many presently studied qubit platforms is a quite costly and time-consuming task [45], they also come with important benefits. On the one hand, under fairly general conditions, it is impossible to target certain states, e.g., Green-Horne-Zeilinger (GHZ) states, by driven-dissipative or passive steering protocols [46], while they are accessible to active steering [40]. On the other hand, QEC is limited to stabilizer codes, while active steering could eventually also work for non-stabilizer codes.
@@
In Sec. III, we present numerical results obtained from the SME. We show that for both $N=1$ and $N=2$, a sharp threshold occurs at a (non-universal) error rate $\Gamma_{c}$. The precise value of $\Gamma_{c}$ depends on the balance between steering processes subject to active decision making and measurement- as well as error-induced stochastic relaxation processes. Our simulation results show that the threshold behavior is clearly visible in the state fidelity and the purity. The threshold is characterized by a purity gap closing as a function of $\Gamma/\Delta$.
@@ II Active steering in the presence of noise
Figure 1: Illustration of the active steering scheme for (a) $N=1$ and (b) $N=2$ system qubits. For $N=1$, the system (“s”, orange square) qubit is initialized at time $t=0$ in the state $|0\rangle$, and the detector (“d”, blue circle) is initialized in $|0\rangle_{d}$. After the time step $\delta t$, during which the Hamiltonian $H$ in Eq. (1) entangles the system and detector qubits (indicated by the dashed line connected system and detector qubits), one performs a projective measurement of the detector in the computational basis. Based on the measured outcome, the averaged expected changes $\overline{(dO_{s})_{\alpha}}$ in Eq. (11) of a system operator $O_{s}$ are calculated for all possible “steering operators” in Eq. (10), with $\alpha=1,\ldots,7$, by a classical computation (“control unit”). We allow for error channels acting on the system qubit during the protocol. The observable $O_{s}$ is chosen such that the eigenstate for its maximal eigenvalue is the desired target state $|\Psi_{f}\rangle$. Maximizing $\overline{(dO_{s})_{\alpha}}$ over the set $\{\alpha\}$ then dictates the steering operator (and thus $H$) selected during the subsequent time step. The detector is always re-initialized in the state $|0\rangle_{d}$. This scheme is iterated until reaching the designated target state $\rho_{f}=|\Psi_{f}\rangle\langle\Psi_{f}|$ with high fidelity and high purity (for sufficiently small error rates). For the $N=2$ case in panel (b), there are no direct couplings between system qubits or between detector qubits. Entanglement between system qubits is generated through entanglement swapping, using Bell pair measurements of the two detector qubits [40]. We again allow for local error channels. For details, see main text.
@@ II.1.1 General formulation
where $H_{s}$ ($H_{d}$) acts only on the system (detector) qubit and $H_{sd}$ is a system-detector coupling. We write $H_{s}={\bf s}\cdot\bm{\sigma}$ with $\bm{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ and a real vector ${\bf s}$. Similarly, $H_{d}={\bf d}\cdot\bm{\tau}$ with $\bm{\tau}=(\tau^{x},\tau^{y},\tau^{z})$. For $H_{sd}$, we use two-qubit Pauli operators or sums thereof, see below for examples. On top of the unitary time evolution induced by $H$ in Eq. (1), we will later on also include error channels. However, for the moment, we proceed by neglecting errors.
@@
One then re-initializes the detector qubit in the state $|0\rangle_{d}$, such that the total initial state for the next time step is $|\Psi(t+\delta t)\rangle_{\xi}\otimes|0\rangle_{d}$, and iterates the scheme. An important part of the active steering protocol is the choice of the Hamiltonian $H$ in different time steps, see below.
@@
We note that Eq. (II.1.1) tacitly assumes that the measurement time required for performing projective detector measurements is short compared to $\delta t$. This assumption may or may not be valid in practice, depending on the platform. Below we give concrete numbers for the case of an Andreev qubit. However, one could refine our analysis by including an extra time interval after each time step, where the measurement dynamics is explicitly considered and the system qubit remains subject to decoherence and dephasing. We leave such generalizations to future work.
Our active steering protocol aims at maximizing a suitable system observable $O_{s}$ such that the target state is approached over the course of the protocol. For instance, if we wish to prepare the state $|\Psi_{f}\rangle=|0\rangle$, the observable is chosen as $O_{s}=\sigma^{z}$ since the maximal expectation value $\langle O_{s}\rangle=+1$ is only achieved for the desired target state. While we only assume pure initial states throughout the paper, it would be straightforward to extend the discussion to an arbitrary initial state, including mixed states. Note, that in the case of unitary measurement jump operators one will not be able to purify the system’s state, as discussed below. Taking a fixed system-detector coupling $H_{sd}$, we assume that, depending on the previous measurement record, one can select one out of seven steering operators $H_{s,\alpha}$ (with $\alpha=1,\ldots,7$) acting on the system qubit during the following time step. In general, for a given time step, the system Hamiltonian $H_{s}$ then contains a piece $H_{s}^{(0)}$, which is always present, and a steering operator $H_{s,\alpha}$ (when choosing the system Hamiltonian for the next step as done here, rather than the system-detector coupling, the active decision making sometimes is referred to as “active shaking” rather than “active steering”, cf. [42], where shake-and-steer protocols are discussed),
@@
$\overline{\langle dO_{s}\rangle_{\alpha}}=\sum_{\xi=0,1}P^{(\alpha)}_{\xi}\mathrm{Tr}\left(d\rho^{(\alpha)}_{\xi}O_{s}\right),$ (11)
where $\overline{O}$ indicates an average over measurement outcomes. Here, $d\rho_{\xi}^{(\alpha)}$ follows from Eq. (II.1.1) with $H_{s}=H_{s}^{(0)}+H_{s,\alpha}$, and the probabilities $P_{\xi}^{(\alpha)}$ are given in Eq. (9). For the next time step, we then select the steering operator (indexed by $\alpha$) which maximizes Eq. (11). In this manner, on average, we effectively also maximize $\langle O_{s}\rangle$ during the course of the protocol.
@@ II.1.2 Application: Andreev qubit
On the other hand, we assume a negligibly short readout time below. While this assumption is rather unrealistic for transmon qubits, where $100$ ns currently sets the optimal value [55], great progress has been achieved for fast ($<10$ ns) dispersive readout of semiconducting quantum dot qubits [81]. In any case, the time step $\delta t$ of the protocol should be chosen large compared to the measurement time of the detector.
@@ II.2 SME for two system qubits
$\overline{\langle dO_{s}\rangle_{\alpha}}=\sum_{\xi,\eta}P_{\xi,\eta}^{(\alpha)}\,{\rm Tr}\left(d\rho^{(\alpha)}_{\xi,\eta}\,O_{s}\right)$ (23)
@@ II.3 Numerical solution of the SME
An individual measurement-resolved trajectory $\rho(t)$ solving the SME is thereby expressed in terms of an average over $n_{w}\gg 1$ pure-state trajectories $|\Psi(t)\rangle$ solving the SSE (II.3) for different realizations of the stochastic bath variables $x_{\gamma}(t)$ (but the same measurement outcomes). The average over $\{x_{\gamma}(t)\}$ mimics the effect of the corresponding dissipative contributions in the SME. In Sec. III, we choose $n_{w}=100$. This value is sufficiently large to ensure convergence to the limit described by the SME without slowing down the numerical simulation too much. On top of the bath average, we may also average over measurement trajectories in order to compute the averaged state $\overline{\rho(t)}$ and/or averaged physical quantities.
@@ III Error threshold
In this section, we present numerical simulation results for the active steering protocols in Sec. II. In Sec. III.1, we study the case of a single Andreev qubit, see Sec. II.1.2, followed by the two-qubit case in Sec. III.2. For clarity, for $N=2$, we assume identical parameters for both qubit-detector subsystems and also identical steering couplings ($J_{1}=J_{2}=J$). In general, we find that our protocol is efficient if $J$ and the system-detector coupling strength $\Lambda$ are of comparable magnitude. The detailed choice of these parameters is then not important, ensuring the robustness of the protocol. As observed from the simulation results below, the number of time steps needed for approaching the final state is $\sim(J\delta t)^{-1}$. Below, we either consider equal error rates for amplitude and phase noise, $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$, or we study the case of only phase noise with $\Gamma_{\rm AD}=0$.
As shown below, for both $N=1$ and $N=2$, there is a sharply defined threshold in the error rate $\Gamma$ which separates two distinct regimes. This threshold, $\Gamma=\Gamma_{c}$, is characterized by a purity gap closing. Below the threshold, $\Gamma<\Gamma_{c}$, autonomous state stabilization is successful and yields a state of high purity, equipped with high fidelity to the target state. Comparing the cases $N=1$ and $N=2$, both these properties become sharper for $N=2$, i.e., both the purity and the fidelity of the ensemble-averaged state remain closer to their ideal values upon increasing $\Gamma$. This behavior suggests that for larger $N$, one approaches a true phase transition like in QEC. However, with the present approach, we cannot address this exciting question. Beyond the threshold, $\Gamma>\Gamma_{c}$, autonomous state stabilization fails. The average state then converges to the one chosen by the remaining generator of dynamics, instead of the desired target state. The final state can be a pure dark state (for combined phase and amplitude noise) or an infinite-temperature state (for dephasing alone).
@@ III.1 One system qubit
Representative numerical simulation results for the time evolution are illustrated in Fig. 3, where we show the system state $\overline{\rho(t)}$ averaged over $1000$ measurement trajectories. Using the Bloch representation, we write $\overline{\rho(t)}=\frac{1}{2}(\sigma^{0}+{\bf r(t)}\cdot\bm{\sigma})$ with the identity $\sigma^{0}$. In Fig. 3, we track the vector ${\bf r}(t)$ in the Bloch unit ball, $|{\bf r}(t)|\leq 1$, at different time steps. For all panels in Fig. 3, we start from $|\Psi(0)\rangle=|+\rangle$ and consider the target state $|\Psi_{f}\rangle=|0\rangle$, corresponding to $O_{s}=\sigma_{z}$ for the active decision making scheme in Sec. II.1.
@@
Figure 4: Time-dependent fidelities, see Eq. (27), obtained from an individual measurement trajectory for $N=2$ (main panel) and $N=1$ (inset) system qubits under the active steering protocol with error rate $\Gamma/\Delta=0.001$, where $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. We use the same parameters as in Fig. 3 but with $\Delta\delta t=0.05$. For $N=2$, both system-plus-detector subsystems in Fig. 1 have identical couplings $\Lambda/\Delta=0.49$. The initial system state is $|\Psi(0)\rangle=|++\rangle$ for $N=2$ ($|\Psi(0)\rangle=|+\rangle$ for $N=1$), and the target state is the symmetric even-parity Bell state $|\Phi_{0,+}\rangle$ in Eq. (25) for $N=2$ ($|\Psi_{f}\rangle=|0\rangle$ for $N=1$). Main panel: Fidelity in Eq. (27) vs number of time steps for $N=2$. We show the fidelities with respect to each of the four Bell states. Inset: Fidelity vs number of times steps for $N=1$ with respect to the states $|0\rangle$ and $|1\rangle$.
@@
This real number is bounded by $0\leq F(t)\leq 1$, with $F=1$ only if $\rho(t)=|\Psi_{f}\rangle\langle\Psi_{f}|$ is fully converged. When averaged over many measurement trajectories, we denote the averaged fidelity by $\overline{F_{|\Psi_{f}\rangle}(t)}$. The inset of Fig. 4 shows the time-dependent fidelities $F_{|0\rangle}(t)$ and $F_{|1\rangle}(t)$ for an individual measurement trajectory with $|\Psi(0)\rangle=|+\rangle$ and $|\Psi_{f}\rangle=|0\rangle$, where the active steering protocol is subject to very weak amplitude and phase damping, $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. We observe that apart from occasional and inevitable quantum jump events, the state is reliably steered towards the desired target state $|0\rangle$. If a quantum jump (with $\xi=1$) occurs, it comes with a large change in the fidelity. However, the protocol “repairs” this glitch, resulting from an unlikely and unfavorable measurement outcome, quickly by itself again. Our active steering protocol can thus be viewed as an implementation of autonomous state stabilization.
Figure 5: Error rate dependence of the fidelity reached at late times of the $N=1$ steering protocol with $|\Psi(0)\rangle=|+\rangle$ and $|\Psi_{f}\rangle=|0\rangle$. Both amplitude and phase noise have been included with $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. Note the logarithmic scales for $\Gamma/\Delta$. We use the same parameters as in Fig. 3. Main panel: Fidelity $\overline{F_{|0\rangle}}$ vs $\Gamma/\Delta$, averaged over $500$ measurement trajectories, for several $\Delta\delta t$. The threshold, $\Gamma_{c}/\Delta\simeq 0.1$, is marked by an arrow. Upper right inset: Averaged fidelity (red curve) and variance (bars) vs $\Gamma/\Delta$ for $\Delta\delta t=0.03$. The variance has been computed from 500 trajectories. Lower left inset: Same as main panel but for $\Delta\delta t=0.03$ and several $J/\Delta$.
@@
For the same $N=1$ protocol as just discussed, we next turn to the asymptotic long-time value of the averaged fidelity, $\overline{F_{|0\rangle}}$, as a function of the error rate $\Gamma$. We again consider the case where both amplitude and phase noise are present. In general, $\overline{F_{|0\rangle}}$ depends on $\delta t$ for otherwise fixed parameters. (Using $\Delta$ as reference energy, the dependence on $\delta t$ is encoded by the dimensionless parameter $\Delta\delta t$.) However, we observe from the main panel of Fig. 5 that results for different $\delta t$ (approximately) fall onto a single scaling curve when plotting the fidelity as a function of $\Gamma/\Delta$. The scaling curve is characterized by a sharp threshold error rate $\Gamma_{c}$ where a qualitative change happens. For the present parameters, we find $\Gamma_{c}/\Delta\simeq 0.1$. For a fixed value of $\delta t$, the lower left inset of Fig. 5 shows how the curves $\overline{F}$ vs $\Gamma/\Delta$ change when $J/\Delta$ is varied. Remarkably, both the threshold value $\Gamma_{c}$ and the scaling property are basically unaffected by changing $J$ as long as we are in the regime $J\gtrsim\Lambda$, where steering couplings are significant and the protocol remains efficient. However, the scaling feature of the fidelity now breaks down for strong damping, $\Gamma>\Gamma_{c}$, where the precise value of $J/\Delta$ matters.
@@
For $\Gamma\ll\Gamma_{c}$, we conclude that amplitude and phase noise do not harm the active steering protocol. One can still achieve high fidelities, and it is possible to approach and/or stabilize the desired target state with high accuracy. However, with increasing error rate, upon approaching $\Gamma=\Gamma_{c}$, a qualitative and sudden change in the average fidelity occurs. Apart from the fidelity, this threshold also shows up in other diagnostic measures as detailed below. The upper right inset of Fig. 5 shows the variance of the fidelity (represented by error bars). For $\Gamma\ll\Gamma_{c}$, the variance is small since basically only the random measurement outcomes cause fidelity variations between different measurement trajectories. As $\Gamma$ grows, however, also the variance increases since stochastic bath forces now become significant and cause a larger spread.
Figure 6: Purity $P$ in Eq. (28) vs $\Gamma/\Delta$ for $N=1$ (main panel) and $N=2$ (inset), in the presence of amplitude and phase damping, $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. Note the logarithmic scales for $\Gamma/\Delta$. We use the same parameters as well as initial and target states as in Fig. 4, with $\Delta\delta t=0.03,\Lambda/\Delta=0.98$ (for $N=1$), and $\Delta\delta t=0.03,\Lambda/\Delta=0.49$ (for $N=2$). Averages are over $500$ measurement trajectories.
@@
Next, in the main panel of Fig. 6, we show the purity $P$ of the late-time averaged system state,

### v2

@@ I Introduction
Below we explore a route to state preparation and stabilization, which builds upon the recently developed “active steering” protocols [35, 36, 37, 38, 39, 40, 41]. (We here follow Ref. [42] and use the notion “steering” as proxy for “guiding” the quantum system; this notion differs from “quantum steering” in quantum information theory [43].) In active steering protocols, one employs feedback operations based on weak measurements of the system, instead of performing projective syndrome measurements as in standard QEC. Typically, in weak measurements, the system qubits of interest are weakly coupled to detector qubits which are then projectively measured [19]. The key idea is then to perform periodically repeated, non-invasive measurements of the system state, where the detector output determines the application of subsequent operations in each time step, e.g., in order to prepare, stabilize, or manipulate a desired quantum state. In the broader context of quantum control, active steering protocols are examples of state-based feedback protocols and measurement-based feedback control. This type of quantum control has a long history in quantum information and quantum optics [19], with one of the main applications being continuous-time error correction, see, e.g., Ref. [44] and references therein. Active steering protocols should be contrasted with “passive steering” protocols [42, 45], where the weak measurement outcomes are simply discarded. In effect, passive steering protocols are equivalent to the driven-dissipative passive protection schemes mentioned above.
While active steering protocols require one to perform many measurements, which in many presently studied qubit platforms is a quite costly and time-consuming task [46], they also come with important benefits. On the one hand, under fairly general conditions, it is impossible to target certain states, e.g., Green-Horne-Zeilinger (GHZ) states, by driven-dissipative or passive steering protocols [47], while they are accessible to active steering [40]. We note that for stabilizer states like GHZ states, simpler state preparation schemes are available, see, e.g., Refs. [48, 49, 50]. The results of Ref. [47] do not apply to such routes. Importantly, active steering could also work for non-stabilizer states.
@@
In Sec. III, we present numerical results obtained from the SME. We show that for both $N=1$ and $N=2$, a sharp threshold occurs at a (non-universal) error rate $\Gamma_{c}$. The precise value of $\Gamma_{c}$ depends on the balance between steering processes subject to active decision making and measurement- as well as error-induced stochastic relaxation processes. Our simulation results show that the threshold behavior is clearly visible in the state fidelity and the purity. The threshold is characterized by a purity gap closing as a function of $\Gamma/\Delta$, where $\Delta$ is the superconducting pairing gap.
@@ II Active steering in the presence of noise
Figure 1: Illustration of the active steering scheme for (a) $N=1$ and (b) $N=2$ system qubits. For $N=1$, the system (“s”, orange square) qubit is initialized at time $t=0$ in the state $|0\rangle$, and the detector (“d”, blue circle) is initialized in $|0\rangle_{d}$. After the time step $\delta t$, during which the Hamiltonian $H$ in Eq. (1) entangles the system and detector qubits (indicated by the dashed line connected system and detector qubits), one performs a projective measurement of the detector in the computational basis. Based on the measured outcome, the averaged expected changes $\langle(dO_{s})_{\alpha}\rangle_{\rm ms}$
in Eq. (11) of a system operator $O_{s}$ are calculated for all possible “steering operators” in Eq. (10), with $\alpha=1,\ldots,7$, by a classical computation (“control unit”). We allow for error channels acting on the system qubit during the protocol. The observable $O_{s}$ is chosen such that the eigenstate for its maximal eigenvalue is the desired target state $|\Psi_{f}\rangle$. Maximizing $\langle(dO_{s})_{\alpha}\rangle_{\rm ms}$ over the set $\{\alpha\}$ then dictates the steering operator (and thus $H$) selected during the subsequent time step. The detector is always re-initialized in the state $|0\rangle_{d}$. This scheme is iterated until reaching the designated target state $\rho_{f}=|\Psi_{f}\rangle\langle\Psi_{f}|$ with high fidelity and high purity (for sufficiently small error rates). For the $N=2$ case in panel (b), there are no direct couplings between system qubits or between detector qubits. Entanglement between system qubits is generated through entanglement swapping, using Bell pair measurements of the two detector qubits [40]. We again allow for local error channels. For details, see main text.
@@ II.1.1 General formulation
where $H_{s}$ ($H_{d}$) acts only on the system (detector) qubit and $H_{sd}$ is a system-detector coupling. We write $H_{s}={\bf s}\cdot\bm{\sigma}$ with $\bm{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ and a real vector ${\bf s}$. Similarly, $H_{d}={\bf d}\cdot\bm{\tau}$ with $\bm{\tau}=(\tau^{x},\tau^{y},\tau^{z})$. For $H_{sd}$, we use two-qubit Pauli operators or sums thereof. On top of the unitary time evolution induced by $H$ in Eq. (1), we will later on, see Eq. (II.1.1) below, also include error channels. However, for the moment, we proceed by neglecting errors.
@@
One then re-initializes the detector qubit in the state $|0\rangle_{d}$, such that the total initial state for the next time step is $|\Psi(t+\delta t)\rangle_{\xi}\otimes|0\rangle_{d}$, and iterates the scheme. An important part of the active steering protocol is the choice of the Hamiltonian $H$ in different time steps.
@@
We note that Eq. (II.1.1) tacitly assumes that the measurement time required for performing projective detector measurements is short compared to $\delta t$. This assumption may or may not be valid in practice, depending on the platform. Below we give concrete numbers for the case of an Andreev qubit. However, one could refine our analysis by including an extra time interval after each time step, where the measurement dynamics is explicitly considered and the system qubit remains subject to decoherence and dephasing. We leave such generalizations to future work. In addition, we neglect the reset time for the detector.
Our active steering protocol aims at maximizing a suitable system observable $O_{s}$ such that the target state is approached over the course of the protocol. For instance, if we wish to prepare the state $|\Psi_{f}\rangle=|0\rangle$, the observable is chosen as $O_{s}=\sigma^{z}$ since the maximal expectation value $\langle O_{s}\rangle=+1$ is only achieved for the desired target state. While we only assume pure initial states throughout the paper, it would be straightforward to extend the discussion to an arbitrary initial state, including mixed states. Taking a fixed system-detector coupling $H_{sd}$, we assume that, depending on the previous measurement record, one can select one out of seven steering operators $H_{s,\alpha}$ (with $\alpha=1,\ldots,7$) acting on the system qubit during the following time step. In general, for a given time step, the system Hamiltonian $H_{s}$ then contains a piece $H_{s}^{(0)}$, which is always present, and a steering operator $H_{s,\alpha}$ (when choosing the system Hamiltonian for the next step as done here, rather than the system-detector coupling, the active decision making sometimes is referred to as “active shaking” rather than “active steering”, cf. Ref. [42], where shake-and-steer protocols are discussed),
@@
$\langle(dO_{s})_{\alpha}\rangle_{\rm ms}=\sum_{\xi=0,1}P^{(\alpha)}_{\xi}\mathrm{Tr}\left(d\rho^{(\alpha)}_{\xi}O_{s}\right),$ (11)
where $\langle O\rangle_{\rm ms}$ indicates an average over measurement outcomes. Here, $d\rho_{\xi}^{(\alpha)}$ follows from Eq. (II.1.1) with $H_{s}=H_{s}^{(0)}+H_{s,\alpha}$, and the probabilities $P_{\xi}^{(\alpha)}$ are given in Eq. (9). For the next time step, we then select the steering operator (indexed by $\alpha$) which maximizes Eq. (11). In this manner, on average, we effectively also maximize $\langle O_{s}\rangle$ during the course of the protocol.
@@ II.1.2 Application: Andreev qubit
On the other hand, we assume negligibly short detector readout and detector reset time scales below. (In case one wants to include the effects of those times explicitly, one has to add the uncoupled but noisy evolution of the system qubits during the measurement and reset periods in the SME.) While this assumption is rather unrealistic for transmon qubits, where a readout time $\sim 100$ ns currently sets the optimal value [59], great progress has been achieved for fast ($<10$ ns) dispersive readout of semiconducting quantum dot qubits [85]. Detector reset times are typically much faster [86]. We note that the time step $\delta t$ of the protocol should be chosen large compared to the measurement (and reset) time of the detector. In fact, the dimensionless parameter $\Delta\delta t$ is limited from both below and from above. The lower limit arises from the measurement and reset) time and, in numerical computations as presented below, from the fact that otherwise the numerical run-times necessary for converging the protocol towards the steady state become very long. The upper limit instead is imposed by the weak measurement limit assumed in our formulation of the theory. For these reasons, we explore only a relatively narrow window of values for $\Delta\delta t$ in Sec. III below. However, already for those values, we find indications for “universal” error thresholds.
@@ II.2 SME for two system qubits
$\langle(dO_{s})_{\alpha}\rangle_{\rm ms}=\sum_{\xi,\eta}P_{\xi,\eta}^{(\alpha)}\,{\rm Tr}\left(d\rho^{(\alpha)}_{\xi,\eta}\,O_{s}\right)$ (23)
@@ II.3 Numerical solution of the SME
An individual measurement-resolved trajectory $\rho(t)$ solving the SME is thereby expressed in terms of an average over $n_{w}\gg 1$ pure-state trajectories $|\Psi(t)\rangle$ solving the SSE (II.3) for different realizations of the stochastic bath variables $x_{\gamma}(t)$ (but the same measurement outcomes). The average over $\{x_{\gamma}(t)\}$ mimics the effect of the corresponding dissipative contributions in the SME. In Sec. III, we choose $n_{w}=100$. This value is sufficiently large to ensure convergence to the limit described by the SME without slowing down the numerical simulation too much. We have checked that this value also suffices to describe the largest damping rates discussed in Sec. III. For a detailed discussion of the convergence properties with respect to $n_{w}$, see Ref. [94]. On top of the bath average, we may also average over measurement trajectories in order to compute the averaged state $\overline{\rho(t)}$ and/or averaged physical quantities.
@@ III Error threshold
In this section, we present numerical simulation results for the active steering protocols in Sec. II. In Sec. III.1, we study the case of a single Andreev qubit, see Sec. II.1.2, followed by the two-qubit case in Sec. III.2. For clarity, for $N=2$, we assume identical parameters for both qubit-detector subsystems and also identical steering couplings ($J_{1}=J_{2}=J$). In general, we find that our protocol is most efficient and robust if $J$ and the system-detector coupling strength $\Lambda$ are of comparable magnitude, not differing by more than one order of magnitude. Within these bounds, the detailed choice of these parameters is not important, ensuring the robustness of the protocol. One can rationalize the robustness of the scheme for $J\approx\Lambda$ by noting that the unitary active feedback operations will be inefficient for small $J$ while measurement-based feedback operations become inefficient for small $\Lambda$. A good compromise is reached for $J\approx\Lambda$. Below, we either consider equal error rates for amplitude and phase noise, $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$, or we study the case of only phase noise with $\Gamma_{\rm AD}=0$.
@@ III.1 One system qubit
Representative numerical simulation results for the time evolution are illustrated in Fig. 3, where we show the system state $\overline{\rho(t)}$ averaged over $1000$ measurement trajectories. Using the Bloch representation, we write $\overline{\rho(t)}=\frac{1}{2}(\sigma^{0}+{\bf r(t)}\cdot\bm{\sigma})$ with the identity $\sigma^{0}$. In Fig. 3, we track the vector ${\bf r}(t)$ in the Bloch unit ball, $|{\bf r}(t)|\leq 1$, at different time steps. For all panels in Fig. 3, we start from $|\Psi(0)\rangle=|+\rangle$ and consider the target state $|\Psi_{f}\rangle=|0\rangle$, corresponding to $O_{s}=\sigma_{z}$ for the active decision making scheme in Sec. II.1. Note that this target state represents an excited state of the Andreev qubit Hamiltonian (12). In fact, we find that the active steering protocol approaches an arbitrary predesignated target state with similar convergence rate.
@@
Figure 4: Time-dependent fidelities, see Eq. (27), obtained from an individual measurement trajectory for $N=1$ [panel (a)] and $N=2$ [panel (b)] system qubits under the active steering protocol with error rate $\Gamma/\Delta=0.001$, where $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. We use the same parameters as in Fig. 3 but with $\Delta\delta t=0.05$. Panel (a): Fidelity vs number of times steps for $N=1$ with respect to the states $|0\rangle$ and $|1\rangle$, with target state $|\Psi_{f}\rangle=|0\rangle$. Panel (b): Same as in panel (a) but for $N=2$. Both system-plus-detector subsystems in Fig. 1 have identical couplings $\Lambda/\Delta=0.49$. The initial system state is $|\Psi(0)\rangle=|++\rangle$ and the target state is the symmetric even-parity Bell state $|\Phi_{0,+}\rangle$ in Eq. (25). We show the fidelities with respect to each of the four Bell states.
Figure 5: Error rate dependence of the fidelity reached at late times of the $N=1$ steering protocol with $|\Psi(0)\rangle=|+\rangle$ and $|\Psi_{f}\rangle=|0\rangle$. Both amplitude and phase noise have been included with $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. Note the logarithmic scales for $\Gamma/\Delta$. We use the same parameters as in Fig. 3. Panel (a): Fidelity $\overline{F_{|0\rangle}}$ vs $\Gamma/\Delta$, averaged over $500$ measurement trajectories, for several values of $\Delta\delta t$. The threshold, $\Gamma_{c}/\Delta\simeq 0.1$, is marked by an arrow. Panel (b): Same as panel (a) but for $\Delta\delta t=0.03$ and several values of $J/\Delta$. Panel (c): Averaged fidelity (red curve) and variance (black bars) vs $\Gamma/\Delta$ for $\Delta\delta t=0.01$. The variance has been computed from 500 trajectories.
Figure 6: Purity $P$ in Eq. (28) vs $\Gamma/\Delta$ for $N=1$, including amplitude and phase damping with $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. Note the logarithmic scales for $\Gamma/\Delta$. We use the same parameters as well as initial and target states as in Fig. 4(a), with $\Delta\delta t=0.03$ and $\Lambda/\Delta=0.98$. Averages are over $500$ measurement trajectories.
@@
This real number is bounded by $0\leq F(t)\leq 1$, with $F=1$ only if $\rho(t)=|\Psi_{f}\rangle\langle\Psi_{f}|$ is fully converged. When averaged over many measurement trajectories, we denote the averaged fidelity by $\overline{F_{|\Psi_{f}\rangle}(t)}$. Figure 4(a) shows the time-dependent fidelities $F_{|0\rangle}(t)$ and $F_{|1\rangle}(t)$ for an individual measurement trajectory with $|\Psi(0)\rangle=|+\rangle$ and $|\Psi_{f}\rangle=|0\rangle$, where the active steering protocol is subject to very weak amplitude and phase damping, $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. We observe that apart from occasional and inevitable quantum jump events, the state is reliably steered towards the desired target state $|0\rangle$. If a quantum jump (with $\xi=1$) occurs, it comes with a large change in the fidelity. However, the protocol “repairs” this glitch, resulting from an unlikely and unfavorable measurement outcome, quickly by itself again. Our active steering protocol can thus be viewed as an implementation of autonomous state stabilization. In general, our simulation results show that the number of time steps needed for approaching the final state is $\sim(J\delta t)^{-1}$.
@@
For the same $N=1$ protocol as just discussed, we next turn to the asymptotic long-time value of the averaged fidelity, $\overline{F_{|0\rangle}}$, as a function of the error rate $\Gamma$. We again consider the case where both amplitude and phase noise are present. In general, $\overline{F_{|0\rangle}}$ depends on $\delta t$ for otherwise fixed parameters. (Using $\Delta$ as reference energy, the dependence on $\delta t$ is encoded by the dimensionless parameter $\Delta\delta t$.) However, we observe from Fig. 5(a) that results for different $\delta t$ (approximately) fall onto a single scaling curve when plotting the fidelity as a function of $\Gamma/\Delta$. The scaling curve is characterized by a sharp threshold error rate $\Gamma_{c}$ where a qualitative change happens. For the present parameters, we find $\Gamma_{c}/\Delta\simeq 0.1$. In order to investigate how this critical value changes when changing physical model parameters for a fixed value of $\Delta\delta t$, Fig. 5(b) shows how the curves $\overline{F}$ vs $\Gamma/\Delta$ change when $J/\Delta$ is varied. Remarkably, both the threshold value $\Gamma_{c}$ and the scaling property are basically unaffected by changing $J$ as long as we are in the regime $J\gtrsim\Lambda$, where steering couplings are significant and the protocol remains efficient. However, the scaling feature of the fidelity now breaks down for strong damping, $\Gamma>\Gamma_{c}$, where the precise value of $J/\Delta$ matters.
@@
For $\Gamma\ll\Gamma_{c}$, we conclude that amplitude and phase noise do not harm the active steering protocol. One can still achieve high fidelities, and it is possible to approach and/or stabilize the desired target state with high accuracy. However, with increasing error rate, upon approaching $\Gamma=\Gamma_{c}$, a qualitative and sudden change in the average fidelity occurs. Apart from the fidelity, this threshold also shows up in other diagnostic measures as detailed below. We note that Fig. 5(c) shows the variance of the fidelity (represented by error bars). For $\Gamma\ll\Gamma_{c}$, the variance is small since basically only the random measurement outcomes cause fidelity variations between different measurement trajectories. As $\Gamma$ grows, however, also the variance increases since stochastic bath forces now become significant and cause a larger spread.
@@
Next, in Fig. 6, we show the purity $P$ of the late-time averaged system state,